# WikiDBGraph: A Data Management Benchmark Suite for Collaborative Learning over Database Silos thanks: *Equal Contribution

### v2

@@ V-A Experimental Settings
We select the top 2,000 database pairs with the highest similarity scores from WikiDBGraph and, among them, choose two representative pairs that have distinct TIDs yet exhibit high similarity in schema embeddings.
Algorithms
We evaluate representative collaborative learning algorithms spanning horizontal federated learning (HFL), vertical federated learning (VFL), and split learning. These algorithms include:
  • Solo: Training on data from a single database.
  • FedAvg [46]: A standard HFL algorithm in which a server averages client model updates.
  • FedProx [47]: An HFL algorithm to mitigate data heterogeneity ($\mu=0.001$).
  • FedOV [48]: An HFL method designed for data or feature overlap scenarios.
  • SplitNN [10]: A split-learning algorithm for vertically partitioned data, where the model is split across clients.
  • Combined: Centralized training on all clients’ data.
Model and Training Hyperparameters
@@
The base model for all algorithms is a two-layer neural network with [64, 32] hidden dimensions. For SplitNN, this architecture is partitioned such that each client’s local model computes a 64-dimensional embedding, and a central server performs aggregation with the final 32-dimensional layer. All models are trained using the Adam optimizer [49] with a learning rate of $10^{-4}$ and a weight decay of $10^{-5}$.
Evaluation Metrics
@@
We partition the data into an 80% training set and a 20% testing set. To provide a comprehensive performance comparison, we report four standard classification metrics: Accuracy, Precision, Recall, and F1-score. All reported results are the mean values over five independent runs on the combined test sets from both clients to ensure robustness.
Hardware
@@
The experiments are conducted on a server equipped with AMD EPYC 9654 96-Core Processors (1.1 TB system RAM) and four NVIDIA H100 PCIe GPUs (80 GB HBM2e memory each).
Licenses
@@
The WikiDBGraph dataset is released under the Creative Commons Attribution 4.0 International (CC BY 4.0) license. The code of WikiDBGraph is released under the Apache License 2.0.
V-B Automated Data Mining and Evaluation
@@
The first stage, database pair sampling, selects the 2000 pairs with the highest embedding similarity scores from all candidates that satisfy our validation criteria—both database contain at least one table with a minimum of 100 rows. The second stage, label selection, identifies a target column for the learning task. We prioritize classification for simplicity; among all columns with a unique value count between 2 and 50, we select the one with the least number of categories to serve as the label. The third stage, table joining, constructs a comprehensive table for each database by left-joining its tables according to their foreign key relationships, starting from the table containing the selected label. To manage memory, this join process is halted if the table’s instance count exceeds one million. We then remove constant columns and any column with more than 50% missing values, followed by feature standardization. The fourth stage is column alignment, where we harmonize the schemas of the two resulting tables by retaining only their common columns. Column names are matched in a case-insensitive manner after ignoring all non-alphanumeric symbols. Finally, the last stage is training and evaluation. For each client, the aligned data are partitioned into an 80%/20% train/test split. All CL algorithms are trained on the clients’ aligned training sets. To assess improvement, every algorithm—including Solo and Combined—is evaluated on the combined test set from all the clients.
@@ 21 $\mathcal{R}[(D_{A},D_{B})]\leftarrow\{\mathrm{Perf}^{\mathrm{Solo}},\mathrm{Perf}[\cdot]\}$;
V-C Performance of Existing CL Algorithms
@@
The performance distributions of different CL algorithms are shown in Fig. 6, and the absolute improvements over the Solo baseline are shown in Fig. 5. We report only the performance for the 1,207 successful tasks. Two primary observations emerge. First, the average scores of CL algorithms generally lie between the Solo and Combined baselines. Second, performance differences among CL algorithms are small; the notable exception is SCAFFOLD, which is known to be sensitive to hyperparameters and was not carefully tuned here. Both observations align with findings reported in NIIDBench [4], a well-known federated learning benchmark, indicating that evaluation using WikiDBGraph is effective and consistent with prior studies.
@@
Despite the improvement in average performance, the generalization of existing CL algorithms to practical tasks—even when coupled with the automated data-mining pipeline (Alg. 1)—remains limited. Table VII summarizes performance improvements across the 1,207 successful tasks: fewer than half of the tasks benefit from CL. Meanwhile, the distributions in Figure 5 show many cases where collaborative training performed worse than the Solo baseline.
@@
The high rate of non-improvement indicates that the primary bottleneck lies in data preprocessing. Our automated pipeline—which uses simple string-based column alignment—can miss semantic matches or, worse, join dissimilar columns. This flawed preparation yields a “garbage in, garbage out” effect: training on misaligned data degrades collaborative performance below the single-client baseline. Section VI supports this conclusion: manually curated alignments on similar database pairs produce substantial gains, implicating the pipeline rather than the CL algorithms.
@@
While severe data heterogeneity (non-IID)—a known challenge for standard algorithms like FedAvg—is certainly a contributing factor, the stark performance gap between these automated results and the significant gains seen in our manually-curated case studies (Tables VIII, X) points to preprocessing as the more immediate and critical challenge.
@@
Two major limitations of the current preprocessing pipeline are table joining and column alignment. First, the current implementation of the table-joining step is capped at one million rows to manage memory, which can discard valuable data; addressing this likely requires redesigning novel CL algorithms that take databases, rather than prejoined tables, as input. Second, the column-alignment step by comparing string similarity is straightforward but may be suboptimal because it ignores semantic similarity; addressing this calls for a new efficient, semantics-aware column-alignment method to determine which columns are equivalent.
[…]
Figure 5: Performance gain distributions ($\Delta$) of different CL algorithms over the Solo baseline across 2,000 database pairs
@@
(b) Distribution of Precision.
(c) Distribution of Recall.
(d) Distribution of Accuracy.
Figure 6: Final metric distributions achieved by Solo, Combined, and different CL algorithms across 2,000 database pairs
TABLE VII: Summary of automated data mining results with different CL algorithms on 2,000 database pairs
@@
VI Case Study
This section includes three case studies using manually curated data: feature overlap (Section VI-A), instance overlap (Section VI-B), and hybrid overlap (Section VI-C). All cases are novel findings enabled by WikiDBGraph, as their corresponding database pairs have distinct TIDs and no explicit links in Wikidata or WikiDBs.
VI-A Case 1: Two Databases with Feature Overlap
@@
We first evaluate the performance of several neural network (NN) based HFL algorithms, with results summarized in Table VIII. From these results, we draw two primary conclusions. First, all HFL methods consistently and significantly outperform at least one of the Solo baselines. This confirms the effectiveness of WikiDBGraph in finding new databases that enable effective collaborative training. Second, a notable performance gap persists between all HFL algorithms and the centralized training baseline (Combined). This behavior is characteristic of the challenges posed by non-identically and independently distributed (non-i.i.d.) data, a well-known issue in federated learning [4].
To demonstrate broader applicability, we extend our evaluation to tree-based models, known as FedTree [44] (Table IX). The results consistently show that both the Combined and FedTree models outperform their respective Solo baselines. This indicates that the benefits of collaboration extend beyond NN, further validating the effectiveness of WikiDBGraph in identifying valuable opportunities for collaborative learning.
TABLE VIII: Horizontal federated learning accuracy on feature-overlapped databases (six-class classification)
@@
TABLE IX: FedTree accuracy on feature overlap
@@
VI-B Case 2: Two Databases with Instance Overlap
@@
TABLE X: SplitNN accuracy on instance-overlapped databases (49-class classification)
@@
The results of NN-based algorithms, summarized in Table X, yield two key findings. First, the split learning approach significantly outperforms the Solo baseline across all evaluation metrics. This result validates our hypothesis that enriching the feature set by incorporating data from the correlated gene table leads to substantial performance gains. Second, we observe a significant performance gap between SplitNN and Combined—the centralized baseline that represents the performance upper bound. While SplitNN captures the majority of the potential improvement, this gap highlights the inherent challenges of distributed model training compared to a centralized approach. In summary, these findings strongly underscore the practical value of leveraging instance-overlapped databases for collaborative learning evaluation.
Further experiments with a tree-based VFL algorithm, SecureBoost [45], are presented in Table XI. The results reinforce our primary findings: across different algorithms, including specialized ones, extending the feature set consistently yields a significant performance boost. This demonstrates the general applicability of our approach and further validates the effectiveness of WikiDBGraph in identifying valuable pairs for feature enrichment.
TABLE XI: SecureBoost accuracy on instance overlap
@@
VI-C Case 3: Multiple databases with Hybrid Overlap
Figure 7: Hybrid overlap scenario: a connected component of six databases in WikiDBGraph with edge color/boldness indicating similarities.

### v3

@@ VI-A Experimental Settings
Data. We select the top 2,000 database pairs with the highest similarity scores from WikiDBGraph and, among them, choose two representative pairs that have distinct TIDs yet exhibit high similarity in schema embeddings.
Algorithms. We evaluate representative collaborative learning algorithms spanning horizontal federated learning (HFL), vertical federated learning (VFL), and split learning. These algorithms include: Solo: Training on data from a single database. FedProx [28]: An HFL algorithm to mitigate data heterogeneity ($\mu=0.001$). FedOV [29]: An HFL method designed for data or feature overlap scenarios. SplitNN [10]: A split-learning algorithm for vertically partitioned data, where the model is split across clients. SFL [30]: A personalized federated learning method on a graph of clients. Combined: Centralized training on all clients’ data.
@@
Model and Training Hyperparameters. The base model for all algorithms is a two-layer neural network with [64, 32] hidden dimensions. For SplitNN, this architecture is partitioned such that each client’s local model computes a 64-dimensional embedding, and a central server performs aggregation with the final 32-dimensional layer. All models are trained using the Adam optimizer [57] with a learning rate of $10^{-4}$ and a weight decay of $10^{-5}$.
@@
Evaluation Metrics. We partition the data into an 80% training set and a 20% testing set. To provide a comprehensive performance comparison, we report four standard classification metrics: Accuracy, Precision, Recall, and F1-score. All reported results are the mean values over five independent runs on the combined test sets from both clients to ensure robustness.
@@
Hardware. The experiments are conducted on a server equipped with AMD EPYC 9654 96-Core Processors (1.1 TB system RAM) and four NVIDIA H100 PCIe GPUs (80 GB HBM2e memory each).
@@
Licenses. The WikiDBGraph dataset is released under the Creative Commons Attribution 4.0 International (CC BY 4.0) license. The code of WikiDBGraph is released under the Apache License 2.0.
VI-B Evaluation of Benchmark Design
This subsection evaluates the embedding-based relationship predictor against baselines and quantifies the contribution of each benchmark design choice via ablation studies.
Effect of prompt design. Table VI demonstrates the importance of prompting with both schema and data. The combined approach (Full) achieves the best overall performance, indicating that schema definitions and data samples offer complementary context.
Effect of encoder model. Table VII shows that the overall conclusions are consistent across encoder backbones and the string-based schema matching. Contrastive learning substantially improves AUC-ROC by 0.03–0.13 for different encoders. For instance, replacing BGE-M3 with all-mpnet-base-v2 [58] yields comparable performance after contrastive learning, suggesting that the gains primarily come from the proposed contrastive objective rather than a specific pretrained encoder.
Effect of negative samples. Fig. 5(a) shows that varying the number of negative samples has only a marginal impact on AUC. In particular, performance does not improve further beyond the default setting and slightly degrades when substantially increasing the number of negatives. We attribute this to the class imbalance in our data, where overly many negatives may bias the contrastive objective toward easy non-matches.
Effect of sample size in prompt. Fig. 5(b) shows that, once at least one tuple is included, the number of sampled tuples in the training prompt has only a limited impact on performance. Increasing the sample size primarily lengthens the context and raises computation cost, while yielding negligible improvements in AUC.
Effect of similarity threshold. Fig. 6 analyzes how the similarity threshold $\tau$ affects graph construction. As $\tau$ decreases, both the average node degree (Fig. 6(a)) and the total number of edges (Fig. 6(b)) grow rapidly, shifting the graph from sparse (dominated by highly similar pairs) to dense (containing many mid-similarity edges). Since most existing CL methods primarily work on highly similar pairs, we adopt a relatively high $\tau$ in our main setting for methodological compatibility. When more advanced CL techniques become available for mid-similarity regimes, our benchmark can readily support them by evaluating on pairs induced by lower thresholds.
TABLE VI: Performance of relationship prediction (optimal threshold, 5 seeds). Contrast.: contrastive learning. Full: Schema + Data
Configuration Method Accuracy AUC-ROC F1
BGE-M3 Original 0.9016±0.0026 0.9663±0.0012 0.9005±0.0029
Schema Only Contrast. 0.9622±0.0015 0.9908±0.0006 0.9618±0.0014
Data Only Contrast. 0.9766±0.0011 0.9964±0.0004 0.9765±0.0012
BGE-M3 + Full Contrast. 0.9772±0.0007 0.9967±0.0004 0.9770±0.0007
TABLE VII: Different encoder model’s test performance (optimal threshold, 5 seeds). Contrast.: contrastive learning.
Encoder Model Method AUC-ROC Precision Recall
String Match (Jaccard) - 0.7367±0.0051 0.7345±0.0117 0.5661±0.0025
BGE-M3 Original 0.9663±0.0012 0.9107±0.0037 0.8905±0.0066
Contrast. 0.9967±0.0004 0.9851±0.0029 0.9689±0.0027
all-mpnet-base-v2 Original 0.8667±0.0027 0.8369±0.0050 0.7401±0.0044
Contrast. 0.9968±0.0004 0.9846±0.0033 0.9827±0.0044
(a) #Negative samples
(b) #Sample size
Figure 5: Effect of hyperparameters on predictive AUC
(a) Averaged degree
(b) Total #edges
Figure 6: Effect of threshold $\tau$ on the graph properties
VI-C Automated Data Mining and Evaluation
@@
The first stage, database pair sampling, selects the 2,000 pairs with the highest embedding similarity scores from all candidates that satisfy our validation criteria—both databases contain at least one table with a minimum of 100 rows. The second stage, label selection, identifies a target column for the learning task. We prioritize classification for simplicity; among all columns with a unique value count between 2 and 50, we select the one with the least number of categories to serve as the label. The third stage, table joining, constructs a comprehensive table for each database by left-joining its tables according to their foreign key relationships, starting from the table containing the selected label. To manage memory, this join process is halted if the table’s instance count exceeds one million. These heuristics enable a practical “zero-human-in-the-loop” pipeline, where the 2–50 value constraint filters out identifiers and continuous variables to ensure uniform classification tasks.
We then remove constant columns and any column with more than 50% missing values, followed by feature standardization. The fourth stage is column alignment, where we harmonize the schemas of the two resulting tables by retaining only their common columns. Column names are matched in a case-insensitive manner after ignoring all non-alphanumeric symbols. As an optional final quality-control step, we manually inspect each candidate task for semantic relatedness, column-alignment correctness, and label reasonableness. In total, 1,420 of 2,000 tasks pass and are used for training and evaluation. This step does not affect our conclusions, because the rejected tasks are dominated by invalid or non-executable cases already invalidated by the automatic pipeline (e.g., near-constant labels, too few non-constant features after joining/filtering, or runtime failures that prevent CL from training). Finally, the last stage is training and evaluation. For each client, the aligned data are partitioned into an 80%/20% train/test split. All CL algorithms are trained on the clients’ aligned training sets. To assess improvement, every algorithm—including Solo and Combined—is evaluated on the combined test set from all the clients.
In later stages, column alignment can be tackled via schema matching methods (e.g., [59, 60]) that explicitly recover semantic correspondences between columns. While [60] also employs a contrastive learning framework, it focuses on generating fine-grained contextualized column embeddings to resolve feature overlap, distinct from our work learning whole-database embeddings for global structural representation. Additionally, TURL [61] provides semantic table representations that facilitate table augmentation, further benefiting vertical alignment in instance-overlapped settings. Integrating these schema-level signals into future automated CL systems can significantly enhance overall integration performance.
@@ 21 $\mathcal{R}[(D_{A},D_{B})]\leftarrow\{\mathrm{Perf}^{\mathrm{Solo}},\mathrm{Perf}[\cdot]\}$;
VI-D Performance of Existing CL Algorithms
@@
The performance distributions of different CL algorithms are shown in Fig. 9. We report only the performance for the 1,207 that successfully completed training. Two primary observations emerge. First, the average scores of CL algorithms generally lie between the Solo and Combined baselines. Second, performance differences among CL algorithms are small; the notable exception is SCAFFOLD, which is known to be sensitive to hyperparameters and was not carefully tuned here. Both observations align with findings reported in NIIDBench [4], a well-known federated learning benchmark, indicating that evaluation using WikiDBGraph is effective and consistent with prior studies.
@@
Despite the improvement in average performance, the generalization of existing CL algorithms to practical tasks—even when coupled with the automated data-mining pipeline (Alg. 1)—remains limited. Table VIII summarizes performance improvements across the 1,207 successful tasks: fewer than half of the tasks benefit from CL.
@@
The high rate of non-improvement indicates that the primary bottleneck lies in data preprocessing. Our automated pipeline—which uses simple string-based column alignment—can miss semantic matches or, worse, join dissimilar columns. This flawed preparation yields a “garbage in, garbage out” effect: training on misaligned data degrades collaborative performance below the single-client baseline. Section VII supports this conclusion: manually curated alignments on similar database pairs produce substantial gains, implicating the pipeline rather than the CL algorithms.
@@
While severe data heterogeneity (non-IID)—a known challenge for standard algorithms like FedAvg—is certainly a contributing factor, the stark performance gap between these automated results and the significant gains seen in our manually-curated case studies (Tables IX, XI) points to preprocessing as the more immediate and critical challenge.
@@
Two major limitations of the current preprocessing pipeline are table joining and column alignment. First, the current implementation of the table-joining step is capped at one million rows to manage memory, which can discard valuable data; addressing this likely requires redesigning novel CL algorithms that take databases, rather than prejoined tables, as input. Second, the column-alignment step by comparing string similarity is straightforward but may be suboptimal because it ignores semantic similarity; addressing this calls for a new efficient, semantics-aware column-alignment method to determine which columns are equivalent. To validate this, we conduct an additional experiment using DeepJoin [62] for semantics-aware column alignment, while still using BGE-M3 as the base embedding model. As shown in Figure 7, on the same task set (with identical labels), semantic-aware alignment improves the mean F1 score by 0.06–0.09 for FedProx, FedOV, and FedAvg.
Figure 7: F1 improvement of DeepJoin for column alignment
Quantification of Non-IID Heterogeneity. To rigorously compare against synthetic non-IID benchmarks [29], we model the label skew as a multivariate Dirichlet process. We estimate the concentration parameter $\alpha$ using the robust Method of Moments [63], which aligns the theoretical Dirichlet variance with the empirical variance observed across WikiDBGraph pairs. The resulting distribution (Fig. 8) reveals a long-tailed spectrum of heterogeneity. While a substantial portion of pairs exhibit near-IID characteristics ($\alpha>5.0$), we identify a critical cluster of highly heterogeneous data: $24.7\%$ of pairs fall into the high-heterogeneity regime ($\alpha<1.0$), with $18.4\%$ exhibiting extreme non-IID skew ($\alpha<0.5$). This confirms that WikiDBGraph does not merely replicate standard distributions but captures the sharp, “long-tail” non-IID challenges inherent in real-world data silos.
Figure 8: Distribution of estimated Dirichlet-$\alpha$ in WikiDBGraph
@@
(b) Distribution of Accuracy.
Figure 9: Final metric distributions achieved by Solo, Combined, and different CL algorithms
TABLE VIII: Summary of automated data mining results with different CL algorithms on 2,000 database pairs
@@
VII Case Study
This section includes three case studies using manually curated data: feature overlap (Section VII-A), instance overlap (Section VII-B), and hybrid overlap (Section VII-D). All cases are novel findings enabled by WikiDBGraph, as their corresponding database pairs have distinct TIDs and no explicit links in Wikidata or WikiDBs.
VII-A Case 1: Two Databases with Feature Overlap
@@
We first evaluate the performance of several neural network (NN) based HFL algorithms, with results summarized in Table IX. From these results, we draw two primary conclusions. First, all HFL methods consistently and significantly outperform at least one of the Solo baselines. This confirms the effectiveness of WikiDBGraph in finding new databases that enable effective collaborative training. Second, a notable performance gap persists between all HFL algorithms and the centralized training baseline (Combined). This behavior is characteristic of the challenges posed by non-identically and independently distributed (non-i.i.d.) data, a well-known issue in federated learning [4].
To demonstrate broader applicability, we extend our evaluation to tree-based models, known as FedTree [55] (Table X). The results consistently show that both the Combined and FedTree models outperform their respective Solo baselines. This indicates that the benefits of collaboration extend beyond NN, further validating the effectiveness of WikiDBGraph in identifying valuable opportunities for collaborative learning.
TABLE IX: Horizontal federated learning accuracy on feature-overlapped databases (six-class classification)
@@
TABLE X: FedTree accuracy on feature overlap
@@
VII-B Case 2: Two Databases with Instance Overlap
@@
TABLE XI: SplitNN accuracy on instance-overlapped databases (49-class classification)
@@
The results of NN-based algorithms, summarized in Table XI, yield two key findings. First, the split learning approach significantly outperforms the Solo baseline across all evaluation metrics. This result validates our hypothesis that enriching the feature set by incorporating data from the correlated gene table leads to substantial performance gains. Second, we observe a significant performance gap between SplitNN and Combined—the centralized baseline that represents the performance upper bound. While SplitNN captures the majority of the potential improvement, this gap highlights the inherent challenges of distributed model training compared to a centralized approach. In summary, these findings strongly underscore the practical value of leveraging instance-overlapped databases for collaborative learning evaluation.
Further experiments with a tree-based VFL algorithm, SecureBoost [56], are presented in Table XII, reinforcing our primary findings: across different algorithms, including specialized ones, extending the feature set consistently yields a significant performance boost. This demonstrates the general applicability of WikiDBGraph and further validates its effectiveness in identifying valuable pairs for feature enrichment.
TABLE XII: SecureBoost accuracy on instance overlap
@@
VII-C Case 3: Multiple Databases with Instance Overlap
We consider a personalized CL classification task that reflects a policy analytics application. Each client corresponds to an independent country-profile database (DB10626, DB13841, DB17587, DB57240, DB73136) where features overlap only partially. The shared objective is to predict the minimum compulsory education age from locally stored attributes. To ensure the target is well-defined and comparable across databases, we discretize numeric ages into four ordinal classes (<5, 5, 6, 7). These five databases form a five-node, sixteen-edge subgraph, equipped with the WikiDBGraph node and edge properties. We then train a GNN-based CL method (SFL [30]) that consumes these graph properties; the results are reported in Table XIII.
TABLE XIII: SFL experiment results (per-DB averages over 5 seeds). Best in bold, second-best underlined. Gain (%) denotes the percentage of DBs where the method outperforms Solo, averaged over seeds.
Accuracy F1
[…]
Solo (No Federation) 0.8930±0.0636 N/A 0.8708±0.0671 N/A
FedAvg 0.9542±0.0444 20.0% 0.9373±0.0581 20.0%
SFL (No Properties) 0.9358±0.0509 16.0% 0.9138±0.0691 16.0%
SFL (Node Only) 0.9145±0.0756 16.0% 0.8979±0.0838 16.0%
SFL (Edge Only) 0.9559±0.0457 24.0% 0.9407±0.0606 24.0%
SFL (Both) 0.9559±0.0498 24.0% 0.9410±0.0655 24.0%
Table XIII supports two key findings. (1) Node and edge properties are critical for graph-based CL, enabling gains beyond FedAvg: enriching SFL [30] with node and/or edge attributes consistently improves both Accuracy and F1 over the no-property variant and yields the best overall performance, outperforming FedAvg. Notably, edge properties contribute the largest gains—the edge-only configuration nearly matches the full model—suggesting that pairwise cross-database signals are particularly informative for guiding message passing and personalization. (2) Graph-based CL improves over Solo, but structure alone is insufficient: SFL without any graph properties already improves over Solo, indicating that leveraging client connectivity is beneficial; however, it remains slightly below FedAvg, implying that the client graph structure by itself is not enough to beat federated learning baselines without informative node/edge properties.
VII-D Case 4: Multiple databases with Hybrid Overlap